OpenStudy (anonymous):

A fish was caught whose tail weighed 9lb. It's head weighed as much as it's tail plus half it's body. It's body weighed as much as it's head and tail. What did the fish weigh?

OpenStudy (anonymous):

Let t= weight of tail Let b=weight of body Let h= eight of head Given: t= 9 h=t+b/2 (b over 2) b=h+t By substitution: h= 9+b/2 1) 2h=18+b subtracted the b to isolate the 18 2h-b=18 divided the 2 to get 9 b=h+9 2) h-b=-9 subtracted the b to isolate the -9 Then subtract 2h-b=18 from -h+b=9 h=27 Since b = h+9 b=27+9 b=36 The fish weighs 9+27+36=72